The length of a rectangle is four times its width. Its perimeter is 90m. Find its dimensions. Use P= 2l +2w
Answer by checkley77(12844)

L=4W
2L+2W=90
2*4W+2W=90
8W+2W=90
10W=90
W=90/10
W=9 ANS. FOR THE WIDTH.
L=4*9
L=36
PROOF:
2*36+2*9=90
72+18=90
90=90